# A self-supervised framework for learning whole slide representations

### v1

@@ 3.3 S3L Implementation
End-to-end whole slide learning with large batch sizes is infeasible due to the gigapixel scale. To enable efficient training, all unaugmented patch tokens were stored prior to whole slide training, and S3L transformations were applied to patch tokens. Each whole slide was represented as a embedding $\boldsymbol{x}\in\mathbb{R}^{n\times d}$ and a coordinate $\boldsymbol{p}\in\mathbb{Z}^{n\times 2}$. Splitting and masking were implemented as row-wise partitioning and dropout on $\boldsymbol{x},\ \boldsymbol{p}$, respectively. Cropping was implemented as coordinate filtering on $\boldsymbol{p}$. A detailed description of our implementation with pseudocode is in Appendix B.
@@ 4 Experiments
We evaluated our method on two previously benchmarked clinical pathology modalities and tasks: intraoperative brain tumor classification using stimulated Raman histology (SRH) and molecular classification of diffuse gliomas using hematoxylin and eosin (H&E) staining.
S3L was compared to pooling baselines, and to a supervised method, SupCon (Khosla et al., 2020), as the performance upper bound. In addition, we evaluated S3L qualitatively through visualization of attention heatmaps. We evaluated S3L’s robustness and generalizability by training with different patch encoders.
4.1 Datasets and Tasks
Stimulated Raman histology classification
We evaluated S3L on a multiclass histopathological brain tumor diagnosis task using an SRH dataset. SRH is an optical microscopy method that provides rapid imaging of unprocessed biological tissues (Freudiger et al., 2008; Orringer et al., 2017), optimally suited for intraoperative applications. Our SRH dataset was collected at the University of Michigan, and the study has been approved by the Institutional Review Board (HUM00083059). Our imaging protocol follows (Jiang et al., 2022), with informed consent from each patient. The SRH dataset consists of an annotated and an unlabeled set. The annotated set includes 2960 WSIs labeled by board-certified expert pathologists and consists of normal brain tissue and six common brain tumor diagnoses: high-grade glioma, low-grade glioma, meningioma, pituitary adenoma, schwannoma, and metastatic tumor. This annotated set was designed for supervised training and evaluation, and was divided into training and evaluation sets with 2035 and 925 slides, respectively. The unlabeled set encompassing 2919 WSIs was intended for self-supervised pretraining in conjunction with the annotated training set. Detailed slide breakdown for each class can be found in Appendix A.
H&E molecular classification
We also evaluated S3L using publicly available brain tumor H&E stained WSIs from The Cancer Genome Atlas (TCGA) and Digital Brain Tumour Atlas (DBTA) (Roetzer-Pejrimovsky et al., 2022). We focused on the molecular classification of diffuse gliomas, the most common malignant brain tumor. Our dataset consisted of three diffuse glioma subgroups as defined by the World Health Organization (Louis et al., 2021): glioblastoma, astrocytoma, and oligodendroglioma. Molecular classification is a challenging computer vision task because the diagnoses are made via molecular testing (such as genetic sequencing), and are not possible for expert pathologists using H&E images. Our H&E dataset comprises a total of 2650 WSIs from both datasets, with 2309 training slides and 341 slides from the TCGA dataset set for evaluation. H&E Slide composition can be found in Appendix A.
@@ 4.2 Implementation details
We used a ResNet-34 (He et al., 2016) patch encoder to extract 512-dimension features for all patches in each slide. The patch encoder was pre-trained with SimCLR (Chen et al., 2020) and HiDisc (Jiang et al., 2023) for both datasets and was frozen during whole slide training. We used a six-layer transformer (Vaswani et al., 2017) slide encoder with the same dimension as patch tokens and 4 heads, with learnable Fourier features positional embeddings (Li et al., 2021b) for better spatial representation. Learned positional embeddings are visualized in Appendix B.2. We followed common practice for most SSL methods and use a two-layer projection head. For BYOL, we used a one-layer projection head and a one-layer prediction head. We used an AdamW optimizer and a cosine decay scheduler after warm-up in the initial 10% of the iterations. The learn rate was adjusted between $1\times 10^{-3}$ and $1\times 10^{-7}$ to accommodate the unique training dynamics of different SSL methods and tasks. With our slide transformation strategy, we adjusted cropping and masking sizes for each experiment and utilize up to 64 patches per slide for each view. SRH and H&E experiments were conducted over 300 and 400 epochs, respectively, with an effective batch size of up to 1024 WSIs. All experiments were done using one NVIDIA A40 GPU with mixed-precision training. Training with the SimCLR and VICReg losses takes approximately 6 hours and 8 hours for BYOL. SupCon experiments are trained for 100 epochs.
@@ 4.3 Evaluation Protocol
Since linear and finetuning evaluation protocols are sensitive to hyperparameters, such as learning rate, we used a $k$ nearest neighbor ($k$NN) classifier for self-supervised evaluation (Caron et al., 2021; He et al., 2020). For each model, we extracted the whole slide representation for both training and testing data by using all the patches in each slide. The nearest neighbor classifier was then used to match each whole slide representation in the testing dataset to the $k$ most similar whole slide representation in the training dataset based on their cosine similarity. H&E evaluation used only WSI from TCGA training set. We report mean class accuracy (MCA) and F1 scores since both datasets are imbalanced, and standard deviation is computed across three random seeds for each experiment.
@@ 5 Results
We first evaluated the S3L frameworks with different families of SSL objectives. We then compared the generalizability of S3L on different state-of-the-art and out-of-domain patch encoders. We present the ablation studies for S3L and discuss the role of each transformation. Finally, we visualize the self-attention coefficients on the CLS token from the last layer of the whole slide encoder.
5.1 S3L benchmark with different SSL objectives
We benchmark S3L with patch encoder pretrained with SimCLR. Max and mean feature pooling are non-parametric aggregation methods with strong performance for good patch encoders (Ilse et al., 2018; Chen et al., 2022b; Li et al., 2021a; Shao et al., 2021). As shown in Table 1, S3L surpasses max and mean pooling across all self-supervised objectives. Best S3L performance approaches supervised SCM performance upper bound, with an MCA difference of only 2.7% for SRH and 4.0% for H&E. We included an alternative supervised method, CLAM (Lu et al., 2021b), a non-transformer-based SOTA model for whole slide classification. Whole slide SCM transformations with SupCon outperform CLAM by +1.5%. These results demonstrate the flexibility and performance boost of SCM transformations for both SSL and supervised training.
SRH H&E
MCA F1 MCA F1
Max pooling 73.5 75.6 68.2 69.2
Mean pooling 72.5 73.2 68.3 67.9
S3L-SimCLR 77.3 (0.6) 77.6 (0.3) 76.5 (0.5) 76.1 (0.4)
S3L-VICReg 82.3 (0.5) 82.3 (0.3) 72.8 (0.7) 72.4 (0.9)
S3L-BYOL 79.4 (0.4) 78.7 (1.2) 74.6 (0.9) 73.8 (0.9)
CLAM 83.5 (1.3) 82.9 (1.2) 78.9 (2.2) 77.7 (1.8)
Supervised SCM 85.0 (0.5) 84.2 (0.6) 80.5 (1.6) 79.9 (1.4)
Table 1: S3L performance on SRH and H&E self-supervised benchmarks. SimCLR patch tokens are used for training and evaluation. For pooling, mean/max pooling operation is performed over all patch tokens for each slide, and uses kNN evaluation. S3L with SimCLR loss is represented as S3L-SimCLR, for example. Each experiment is run 3 times and a mean (std) is reported.
5.2 Out-of-Distribution generalization of S3L
We explored the generalization ability of S3L with a variety of patch encoders with different model architectures, training algorithms, and dataset domains. A list of SOTA patch encoders is described in Table 2. Based on their training data, patch encoders were grouped into three categories: Out-of-Distribution (OOD), near-domain, and in-domain. ImageNet patch encoder was considered OOD for both SRH and H&E because of the domain gap between WSI and natural images. HIPT (Chen et al., 2022b) was trained with H&E WSIs from TCGA. PLIP (Huang et al., 2023) was trained with images and comments from Twitter. Both include data from multiple organ systems and institutions. Therefore, they were considered near-domain for our brain tumor H&E benchmark and OOD for our SRH benchmark. The SimCLR patch encoder, as in Section 5.1, and HiDisc patch encoder are in-domain since they use the training data from our benchmarks. The performance of S3L with the SimCLR loss is shown in Figure 5, with extended metrics available in Appendix C.
@@ 5.2 Out-of-Distribution generalization of S3L
Training
algorithm
ImageNet ResNet34 ImageNet Cross entropy
HIPT ViT-S/16 All of TCGA DINO
PLIP ViT-B/32 OpenPath CLIP
SimCLR ResNet34 In-domain (see 4.1) SimCLR
HiDisc ResNet34 In-domain (see 4.1) HiDisc
@@
Table 2: Patch encoders. We benchmark S3L with different patch encoders. ImageNet, HIPT, and PLIP patch encoders use off-the-shelf models trained on OOD or near-domain datasets. SimCLR and HiDisc patch encoders are trained with in-domain data.
Figure 5: S3L performance with different patch encoders. For SRH, ImageNet, HIPT, and PLIP patch encoders are pre-trained on OOD data. For H&E, HIPT and PLIP patch encoders are considered near-domain.
S3L compliments patch encoders.
For the SRH benchmark, the patch encoders trained on OOD data fail to generalize, as demonstrated by the poor pooling performance shown in Figure 5. However, the patch encoders learn general visual features that were leveraged by S3L. After training with S3L, WSI representations improved dramatically, as demonstrated by a maximum 18.2% boost in MCA for ImageNet. For the H&E benchmark, S3L compensated for OOD and near-domain patch encoders, surpassing in-domain pooling performance. In addition, S3L closes the performance gap between OOD, near-domain, and in-domain patch encoders, all being within 4% MCA. Overall, these results demonstrate that S3L can enhance whole-slide representation learning using encoders from different domains, thereby reducing the reliance on a specialized in-domain patch encoder. S3L training time ranges from 6-8 GPU hours, while the patch encoder is computationally intensive and it usually takes over 80 GPU hours to train. These results highlight S3L’s versatility and flexibility with different patch encoders and its broader impacts on efficient training for WSI representations.
S3L boosts performance of in-domain patch encoders.
Among the patch encoders, in-domain HiDisc is the highest performing, as demonstrated by the MCA metric for both benchmarks, regardless of the aggregation method. HiDisc patch encoder enforces slide discrimination as an inductive bias during patch training as part of its learning objective. S3L boosts whole slide representations even in the setting of high-quality, in-domain patch encoders. We found that S3L shows more improvement in H&E than the SRH benchmark. This could be due to higher variability between different regions within each H&E WSI. S3L allows the transformer to encode a slide representation from multiple patch tokens, instead of using slide membership of single patches to learn patch representations, as employed in HiDisc.
5.3 Ablation studies
These ablation experiments are conducted under the same configurations unless otherwise noted.
Figure 6: Attention heatmap visualization. Self-attention visualization demonstrates S3L’s ability to distinguish between different tissue phenotypes such as blood, dense tumor, and necrosis. To our knowledge, this is the first self-supervised transformer-based attention maps on full gigapixel WSIs, instead of localized regions.
Transformation choices
Table 3 reports the performance of individual and combinations of slide-level transformations. We leave out splitting alone since it does not reduce the total number of patch tokens for both views, making it prohibitive to train. For the masking and cropping alone, we show that cropping performs better than masking for both SRH and H&E. Combining cropping and masking shows similar performance with cropping alone, but is more efficient to train due to the reduced number of tokens per view. Combining each transformation with splitting improves or maintains performance, showing the benefit of further decreasing mutual information between views. Overall, cropping is the most important transformation, because capturing regional heterogeneity is the most effective and challenging pre-text task.
SRH H&E
MCA F1 MCA F1
M 75.4 (0.9) 76.0 (0.7) 69.9 (0.5) 69.7 (0.5)
C 81.3 (1.5) 81.5 (1.5) 74.3 (0.2) 74.4 (0.2)
S+M 76.1 (0.3) 76.3 (0.8) 69.2 (1.5) 69.3 (1.5)
S+C 82.1 (1.4) 82.2 (1.2) 73.6 (0.6) 73.8 (0.4)
C+M 81.2 (0.5) 81.8 (0.5) 75.5 (2.1) 75.4 (2.2)
S+C+M 82.3 (0.5) 82.3 (0.3) 76.5 (0.5) 76.1 (0.4)
Table 3: Ablation over transformations. S, splitting, C, cropping, M, masking.
Transformation parameters
We further examine the relationship between mutual information and model performance by combining and adjusting the slide transformations. We studied the effect of maximizing the number of tokens included in each view, with SCM and masking only. The results are illustrated in Figure 7. Mutual information levels can be adjusted using SCM, and these transformations should be used in conjunction with each other to optimize model performance. Previous studies have concluded that fewer patch tokens resulted in better model performance (Lazard et al., 2023), but only in the case of using masking transformation alone. Additional transformation ablation studies with crop area sizes are in Appendix C.3.
Figure 7: Max token limit ablation. SCM, splitting-cropping-masking. Shaded area, std across three different random seeds.

### v2

@@ 3.3 SPT Implementation
End-to-end whole slide learning with large batch sizes is infeasible due to the gigapixel scale. To enable efficient training, all unaugmented patch tokens were computed on a frozen patch encoder, and SPT transformations were then applied to patch tokens. Each whole slide was represented as an embedding $\boldsymbol{x}\in\mathbb{R}^{n\times d}$ and a coordinate $\boldsymbol{p}\in\mathbb{Z}^{n\times 2}$. Splitting and masking were implemented as row-wise partitioning and dropout on $\boldsymbol{x},\ \boldsymbol{p}$, respectively. Cropping was implemented as coordinate filtering on $\boldsymbol{p}$. A detailed description of our implementation with pseudocode is in Appendix B.
@@ 4 Experiments
4.1 Benchmarks
We evaluated SPT visual representations on five benchmarks across three clinical tasks, including two different imaging modalities. Additional dataset information and breakdown are in Appendix A.
SRH CNS benchmark.
Stimulated Raman Histology (SRH) is a novel optical microscopy method that enables fast imaging of unprocessed tissues [49, 50]. The benchmark includes six central nervous system (CNS) tumors and normal brains, with 2035 and 925 WSIs for training and evaluation, respectively. These data were collected at University of Michigan, following the imaging protocol in [51], and were labeled by board-certified pathologists. The study has been approved by the Institutional Review Board (HUM00083059), with informed consent from each patient.
H&E glioma molecular classification benchmark.
We also evaluated SPT using publicly available diffuse glioma H&E stained WSIs from the Cancer Genome Atlas (TCGA) and Digital Brain Tumour Atlas (DBTA) [52]. We focused on the classification of three molecular subgroups, as defined by the World Health Organization [53]. Molecular classification is a challenging computer vision task because the diagnoses are made via molecular testing (such as genetic sequencing), and are not possible for expert pathologists using H&E images alone. Our glioma dataset is comprised of 2309 training slides and 341 slides from the TCGA dataset set for evaluation.
TCGA BRCA, TCGA NSCLC, and TCGA RCC benchmarks.
We further evaluated SPT using three widely used TCGA H&E classification benchmarks: 1) invasive breast carcinoma (BRCA) subtyping, 2) non-small cell lung carcinoma (NSCLC) subtyping, and 3) renal cell carcinoma (RCC) subtyping. For these benchmarks, we followed the well-established study design of [5].
@@ 4.2 Implementation details
We trained in-domain ResNet-34 [54] patch encoders to extract features for all patches in each slide. These patch encoders were trained with SimCLR [16] and HiDisc [13]. Additional off-the-shelf and foundational patch encoders are described in Figure 5(a). SPT slide encoders are six-layer transformers [42] with four heads, and learnable Fourier positional embeddings [55] (visualized in Appendix B.2). We used a two-layer projection head for SimCLR and VICReg, and one-layer projection and prediction heads for BYOL. We used AdamW optimizers and cosine decay schedulers after warm-up in the initial 10% of the iterations. The learning rate was adjusted between 10-3 and 10-7 to accommodate the training dynamics of different SSL methods and tasks. With our SPT slide transformation strategy, we adjusted cropping and masking sizes for each experiment and utilized up to 64 patches per slide for each view. We trained ssSPT and suSPT experiments up to 800 and 100 epochs, respectively, with an effective batch size of up to 1024 WSIs. All models were trained with mixed-precision on a NVIDIA A40 GPU, taking up to 8 hours.
@@ 4.3 Evaluation Protocol
We benchmarked SPT using standard linear evaluation protocols. Since linear classifiers are sensitive to hyperparameters [47], we also employed k nearest neighbor (kNN) for direct evaluation. H&E glioma molecular evaluation used only WSI from TCGA. For SRH and H&E Glioma, experiments were repeated with three random seeds, and for TCGA BRCA, NSCLC, and RCC, we used 10-fold cross-validation for error bars, following the protocols in previous work [5, 15]. For the baselines, original embeddings or logits were used when available. We used mean class accuracy (MCA), F1 scores, and area under the receiver operating characteristic (AUC) to evaluate all benchmarks.
@@ 5 Results
We first benchmarked ssSPT performance with self-supervised WSI learning strategies in section 5.1. We compared ssSPT and suSPT with supervised WSI MIL strategies in section 5.2. We then evaluated the ability of SPT to generalize across a wide range of patch encoders in section 5.3. Next, we showed that SPT improves MIL results with SOTA foundation model patch encoders in section 5.4. Finally, we visualized SPT self-attention heatmaps in section 5.5. Full results with error bars, ablation studies on the SPT transformations and parameters, and additional visualizations are in Appendix C.
5.1 ssSPT learns high-quality whole slide representations
We benchmarked ssSPT with self-supervised baselines using in-domain patch encoders. WSI features were evaluated directly using a kNN classifier. As shown in Table 1, ssSPT surpasses all baselines across all self-supervised objectives in all metrics, except for AUC on the highly imbalanced BRCA benchmark. ssSPT outperforms all existing self-supervised methods by a large margin, outperforming the previous best self-supervised method with a 10 and 5 points increase in MCA on the SRH CNS and H&E glioma benchmarks, respectively.
SRH CNS H&E Glioma TCGA BRCA TCGA NSCLC TCGA RCC
MCA F1 AUC MCA F1 AUC MCA F1 AUC MCA F1 AUC MCA F1 AUC
Self-supervised methods
Pooling 72.5 73.2 94.8 68.3 67.9 87.9 58.1 27.4 74.7 75.1 76.5 85.3 84.2 86.3 96.6
HIPT - - - - - - 63.2 39.9 77.5 80.6 80.1 88.9 88.4 89.2 97.4
Giga-SSL 71.2 72.7 94.2 71.6 71.5 89.1 70.4 53.9 85.4 84.3 84.9 92.3 90.0 89.3 97.5
ssSPT (Ours) 82.3 82.3 94.7 76.5 76.1 90.9 72.7 58.3 80.4 86.1 86.1 92.3 91.7 90.5 98.3
Table 1: Self-supervised benchmarks. We use $k$NN classifier to evaluate ssSPT and baselines. We report the best performing SSL objective for ssSPT, with additional SSL objectives in Appendix C Table 11. Mean values are reported here and standard deviations are in Appendix C Table 7.
5.2 SPT outperforms previous fully supervised methods
We benchmarked ssSPT and suSPT with linear evaluation using in-domain patch encoders. As shown in Table 2, SPT outperforms previous fully supervised methods across all tasks. Remarkably, ssSPT outperforms or matches previous fully supervised methods on MCA in SRH CNS (+1.0), H&E glioma (match), TCGA BRCA (+2.2), and TCGA RCC (+1.4) benchmarks. suSPT outperforms existing MIL methods across all five benchmarks on nearly all metrics. In comparison with the best-performing baselines on MCA of these five benchmarks, suSPT achieves a performance increase of 1.2, 0.7, 3.4, 2.1, and 0.3 points, respectively. Thus, SPT provides a performance increase for both self-supervised and supervised learning, generalizing to different tissue types and diagnostic tasks.
SRH CNS H&E Glioma TCGA BRCA TCGA NSCLC TCGA RCC
MCA F1 AUC MCA F1 AUC MCA F1 AUC MCA F1 AUC MCA F1 AUC
Self-supervised methods
Giga-SSL 78.6 75.6 96.1 76.5 73.9 90.7 81.7 62.6 91.2 87.5 87.9 93.7 89.9 87.8 97.7
ssSPT (Ours) 85.4 83.2 97.6 80.0 78.0 91.9 83.0 67.4 89.4 88.0 88.0 94.8 93.4 91.1 98.7
Supervised methods
ABMIL 79.8 78.9 96.0 75.0 74.5 90.2 78.6 62.8 86.1 86.8 86.7 93.7 89.9 89.4 97.8
CLAM 83.5 82.9 96.7 78.9 77.7 88.5 76.8 59.3 85.8 84.3 83.6 92.8 85.6 83.6 97.3
DSMIL - - - - - - 77.0 58.0 83.8 84.9 84.0 92.0 90.1 87.1 97.1
TransMIL 84.4 84.0 84.0 80.0 79.4 91.9 80.8 69.4 90.0 87.4 87.5 95.3 87.4 87.5 95.3
HIPT - - - - - - 75.8 60.2 87.4 88.2 88.2 95.2 92.0 92.1 98.0
suSPT (Ours) 85.6 84.4 97.2 80.7 80.3 93.7 84.2 70.4 90.5 90.3 90.4 95.7 92.3 91.0 98.6
Table 2: SPT benchmarks. We report linear evaluation results on five histology benchmarks. We report the best performing SSL objective for ssSPT, with additional SSL objectives in Appendix C Table 11. Mean values are reported here and standard deviations are in Appendix C Table 8.
5.3 SPT offers performance boost across a range of patch encoders
We explored the generalizability of SPT to different patch encoders using the more challenging H&E Glioma benchmark. As described in Table 5(a), we tested ImageNet, HIPT [5], PLIP [14], and UNI [7], in addition to in-domain patch encoders trained with SimCLR and HiDisc. The ImageNet patch encoder is off-the-shelf and out-of-distribution (OOD). HIPT [5] is near-domain since it was trained with H&E WSIs from TCGA, including multiple organ systems and institutions. PLIP [14] and UNI [7] are visual foundational models for histology, trained with OpenPath and large-scale institutional datasets, respectively. SPT performance is shown in Figure 5(b), with additional results for SRH CNS benchmark in Appendix C Figure 8, and extended metrics with error bars are in Table 9.
For all patch encoders, SPT training significantly improves whole slide representations over pooling baselines. This improvement is the most significant for ImageNet and HIPT patch encoders. As expected, the performance boost is smaller for in-domain patch encoders, especially HiDisc, where slide discrimination was learned during patch training. SPT also achieves a large performance boost using PLIP patch encoders, as it bridges the domain gap between our benchmark and the pre-training dataset OpenPath which is collected on Twitter. As for the state-of-the-art UNI patch encoder, we still observed a large performance boost achieving MCA near 90 points for suSPT. Overall, these results demonstrate that SPT can enhance whole-slide representation learning using encoders from different domains, thereby reducing the reliance on a specialized in-domain patch encoder. SPT training time ranges from 6-8 GPU hours, while the patch encoder is computationally intensive and usually took over 80 GPU hours to train.
@@ 5.3 SPT offers performance boost across a range of patch encoders
ImageNet [54] ResNet34 ImageNet
HIPT [5] ViT-S/16 All of TCGA
SimCLR [16] ResNet34 In-domain (see 4.2)
HiDisc [13] ResNet34 In-domain (see 4.2)
PLIP [14] ViT-B/32 OpenPath
UNI [7] ViT-L/16 Mass-100K + GTEx
@@
(a) Patch encoders. ImageNet is an off-the-shelf model trained on OOD data. HIPT is near-domain for H&E glioma. SimCLR and HiDisc patch encoders are trained with in-domain data. PLIP and UNI are SOTA foundational patch encoders.
(b) SPT results on H&E Glioma.
Figure 5: SPT benchmarks with different patch encoders. ssSPT and suSPT offer performance boosts with a wide range of patch encoders. ssSPT approaches supervised performance upperbound. Additional metrics with error bars are in Appendix C Table 9.
5.4 SPT improves state-of-the-art MIL results
Most recently, foundation models have been developed for computational pathology, promising a general approach for WSI diagnostic tasks. UNI is a foundational patch encoder that achieves state-of-the-art results [7]. The UNI authors found that “ABMIL with UNI features outperforms many sophisticated MIL architectures” [7]. We benchmarked suSPT using UNI patch features, in comparison to ABMIL in Table 3. suSPT outperforms ABMIL in all benchmarks with UNI features, improving upon the previous state-of-the-art performance on these tasks. This showcases that SPT is complementary to the innovative medical foundation models for learning patch representations.
SRH CNS H&E Glioma TCGA BRCA TCGA NSCLC TCGA RCC
MCA F1 AUC MCA F1 AUC MCA F1 AUC MCA F1 AUC MCA F1 AUC
UNI + ABMIL 86.3 86.3 96.3 86.8 85.2 95.5 88.6 79.1 88.8 93.6 90.6 93.4 95.2 93.3 98.5
UNI + suSPT 86.4 86.8 98.4 89.1 88.7 97.3 89.6 82.2 94.7 95.5 95.6 97.6 96.2 94.9 98.9
Table 3: suSPT improves state-of-the-art MIL results with UNI patch features. Mean values are reported here, and standard deviations are in Appendix C Table 10.